# Large Language Model-assisted Autonomous Vehicle Recovery from Immobilization

### v1

@@ Abstract
Despite significant advancements in recent decades, autonomous vehicles (AVs) continue to face challenges in navigating certain traffic scenarios where human drivers excel. In such situations, AVs often become immobilized, disrupting overall traffic flow. Current recovery solutions, such as remote intervention (which is costly and inefficient) and manual takeover (which excludes non-drivers and limits AV accessibility), are inadequate. This paper introduces StuckSolver, a novel Large Language Model (LLM)-driven recovery framework that enables AVs to resolve immobilization scenarios through self-reasoning and/or passenger-guided decision-making. StuckSolver is designed as a plug-in add-on module that operates on top of the AV’s existing perception–planning–control stack, requiring no modification to its internal architecture. Instead, it interfaces with standard sensor data streams to detect immobilization states, interpret environmental context, and generate high-level recovery commands that can be executed by the AV’s native planner. We evaluate StuckSolver on the Bench2Drive benchmark and in custom-designed uncertainty scenarios. Results show that StuckSolver achieves near–state-of-the-art performance through autonomous self-reasoning alone and exhibits further improvements when passenger guidance is incorporated.
@@ I Introduction
After years of development, AVs have progressed beyond controlled environments such as campuses, airports, and ports, or specific public roads with light traffic and simple infrastructure like suburban areas and highways. Today, they are increasingly navigating more complex urban environments. Companies like Baidu and Waymo have deployed large fleets of robotaxis in cities such as Beijing, Wuhan, Phoenix, San Francisco, and Los Angeles, providing public services. For instance, Waymo’s robotaxis offer approximately 150,000 paid rides weekly, collectively traveling over one million miles  [5]. This expansion marks a significant step toward integrating AVs into everyday urban mobility.
@@
Figure 1: Scenario of AV immobilization in traffic. A white AV, carrying non-driving passengers, is traveling in the right lane. When the leading blue vehicle breaks down and the left lane is obstructed by a plastic bag, the AV’s re-planning module fails to generate a new path in the left lane due to the perceived blockage—since it does not recognize that the plastic bag, although detected as an obstacle, is actually traversable. Consequently, the AV comes to a halt and remains stationary indefinitely.
@@
Although AVs can successfully complete most trips, they still face significant challenges in navigating certain traffic scenarios. Factors such as insufficient semantic understanding of surrounding entities and dense interactions at critical spatial locations can cause the decision-making or planning module to fail to generate, or mistakenly filter out, valid solutions, ultimately resulting in immobilization (i.e., becoming stationary or unable to operate normally)  [36, 37]. More critically, when this occurs, they frequently lack effective recovery mechanisms to resume normal operations, as illustrated in Fig. 1. Such incidents can cause substantial traffic disruptions and inconvenience for passengers. For example, a Waymo robotaxi in California became trapped in a roundabout after failing to generate a feasible exit trajectory. The vehicle circled 17 times before being remotely rescued by engineers  [6]. Passenger intervention is another recovery method. For instance, Tesla’s Full Self-Driving (FSD) system exits autonomous mode upon detecting operational limits and prompts the driver to take control [11]. However, these existing methods have notable limitations. Replanning-based approaches essentially re-execute the original decision-making or planning algorithms, but with updated perception inputs that differ from those at the time of immobilization. If the environment has not changed significantly, or if the AV’s understanding of the scene has not improved, the system still fails to generate an effective solution. Remote control requires significant financial and human resources to ensure prompt assistance. Passenger intervention, on the other hand, relies on the occupant’s ability to drive. This leaves elderly, disabled, or non-driving passengers without a viable solution. These challenges underscore the need for more robust and inclusive approaches to enhance AV resilience and accessibility.
@@
Given these limitations, a more efficient and accessible recovery method for AVs is essential. LLMs offer a promising solution due to their extensive knowledge and advanced reasoning capabilities [12]. On one hand, LLMs have the potential to understand complex traffic scenarios and generate logical driving decisions to rescue immobilized AVs, eliminating the need for costly remote control resources [15]. On the other hand, LLMs can process natural language, enabling them to act as a bridge for passenger intervention [16]. By interpreting human commands, LLMs can advise appropriate AV behaviors, making AV rides more accessible and equitable. This capability is particularly beneficial for individuals who do not or cannot drive, ensuring that AV technology serves a broader and more inclusive population.
@@
In this paper, we propose StuckSolver, an LLM-based recovery method designed to address AV immobilization in traffic. StuckSolver is developed as an add-on module that can be seamlessly integrated into existing rule-based AV systems, requiring minimal adaptation. It continuously monitors the surrounding traffic conditions and the AV’s operational status. When immobilization is detected, StuckSolver intervenes by generating recovery decisions and overriding the AV’s original operational plans. In scenarios where StuckSolver is unable to resolve the immobilization autonomously, passengers can provide guidance, with StuckSolver acting as an intermediary to interpret and execute human commands. To evaluate its efficacy, StuckSolver is evaluated on the corner-case-rich benchmark Bench2Drive  [35]. Simulation results demonstrate that StuckSolver effectively recovers immobilized AVs, either by leveraging its own knowledge and reasoning capabilities or by interpreting passenger instructions. These findings highlight the potential of StuckSolver to enhance AV resilience and accessibility in complex traffic scenarios.
@@ II Related work
However, most existing research focuses on isolated tasks—such as perception, decision-making, or planning—letting LLMs operate independently and evaluating their performance within only a single module. While this can demonstrate improvements to specific components of the AV system, it offers little clear evidence of the enhancement in overall driving capability. Some studies have integrated LLMs into a complete AV system [33, 34] and evaluated the driving performance of LLM-powered AVs. However, those efforts primarily focus on normal driving conditions and largely neglect edge cases. Consequently, AVs may struggle in complex or unforeseen scenarios, increasing the risk of prolonged immobilization and compromising both safety and efficiency. Moreover, most of the proposed methods rarely allow passenger intervention, such as providing necessary information or decision guidance. This lack of communication limits the AV’s ability to recover from stuck situations to some extent. These gaps highlight the need for more holistic approaches to fully leverage the potential of LLMs in AV systems, ensuring they can handle complex or unforeseen scenarios scenarios.
Figure 2: An overview of the LLM-assisted AV recovery method. The simulation begins with the AV’s perception module retrieving environmental data from CARLA. The AV system sequentially completes perception, route planning, and decision-making, after which the control module generates steering, throttle, and brake commands. During the simulation, the LLM-based agent, StuckSolver, continuously monitors the AV operations. It retrieves data from the AV system. Using CoT prompting, it performs observation, analysis, and decision-making. If an immobilization scenario is detected, StuckSolver outputs a recovery plan to save the AV. Otherwise, it outputs None, ensuring that normal AV operations remain unaffected. Within the data from AV system, the stationary measures the time the AV remains at a complete stop. The destination flag represents whether the AV reaches its destination. Behavior plan is the list of behaviors. Reason indicates the rationale for the decision. Route replanning? specifies whether route replanning is needed. Route start point defines the start point for replanning.
@@ III LLM-powered Recovery Method
In this section, we present an LLM-powered recovery method designed to enable AVs to autonomously escape from immobilized situations or recover under passenger guidance through natural language commands, eliminating the need for direct driving intervention. Our approach transforms GPT-4o [27] into an agent called StuckSolver by leveraging prompt engineering [28], Chain-of-Thought (CoT) reasoning [29], and OpenAI’s Function Calling API [30] for multimodal information processing, immobilization detection, and multi-step decision-making. StuckSolver operates in a zero-shot mode, requiring no task-specific fine-tuning or additional training. It is integrated into the AV system as an add-on module, continuously monitoring the vehicle’s status during trips. When it detects an immobilization, StuckSolver analyzes the surrounding traffic environment and passenger instructions, then generates a recovery plan that includes a route replanning flag, a route start point, and an action plan to assist the vehicle in regaining normal operation. The overall framework is illustrated in Fig. 2.
@@ III-A StuckSolver
StuckSolver’s role, defined in the system prompt, is to help the AV recover from stuck situations or follow passenger guidance to facilitate recovery. StuckSolver follows designed CoT prompt, sequentially performing environment observation, multi-step analysis, and decision-making, as shown in Fig. 3.
@@
Figure 3: Chain of Thought for StuckSolver. S is an abbreviation for Step. {immobilized?,reason} represents whether the AV is immobilized and the corresponding reason. Here, immobilized? equals 0 if the AV is not immobilized and 1 if the AV is immobilized. The cause indicates the rationale for the decision.
@@
Observation. Observation is the first and most fundamental step in StuckSolver’s reasoning process. A thorough observation phase ensures that the system collects all necessary information before advancing to analysis and decision-making. StuckSolver captures raw images from the vehicle’s front-view camera and extracts semantic information, including traffic control factors and traffic participants. Traffic control factors ($TC$), such as traffic lights state ($TC_{tl}$), traffic signs content ($TC_{ts}$), and work zones ($TC_{wz}$). Traffic participants ($TP$), including vehicles ($TP_{v}$), pedestrians ($TP_{p}$), and obstacles ($TP_{o}$). For $TP_{v}$ and $TP_{p}$, they include type, lane position (ego lane or adjacent lane), and intent. Beyond semantic interpretation, StuckSolver associates the semantic insights with corresponding object measurements from the perception module and enriches it with quantitative attributes such as distance, velocity to facilitate the following analysis.
@@
Analysis. At this stage, the goal is to determine whether the vehicle is stuck and, if so, identify the cause. The analysis process follows two possible branches: with or without passenger guidance. In the absence of passenger guidance (S2.2-2.5), StuckSolver first checks the destination flag to determine whether it has stopped enroute to its destination. If the vehicle is still in motion, StuckSolver outputs None, allowing autonomous operation to continue without intervention. However, if the vehicle has not yet reached its destination and its speed remains below the minimum threshold $v_{min}$ (set to 4.5km/h in this study) for longer than $\delta t$ (set to 1 second in this study), it is flagged as potentially immobilized, triggering further analysis. StuckSolver begins by assessing whether any traffic control elements are influencing the vehicle’s movement. It checks $TC$ to identify possible causes of immobilization. Also, StuckSolver evaluates, based on $TP$, whether the AV has stopped to yield to surrounding traffic participants, or to avoid a potential collision with obstacles. Finally, it provides the analysis results—indicating whether the vehicle is immobilized and the underlying cause—to the decision-making process. When passenger guidance is available (S2.1), StuckSolver interprets the passenger’s intent based on observations and formulates a behavior plan to assist the immobilized vehicle in resuming movement.
@@
Decision-making. If a passenger-suggested behavior plan is available (S3.1), StuckSolver evaluates whether route replanning is necessary to reach the destination. In the absence of passenger guidance (S3.2), StuckSolver assesses the ego state, map information, and available behavior options to generate a safe and traffic-compliant behavior plan. It also determines whether route replanning is required and identifies the appropriate starting point for the new route. Finally, the structured output of StuckSolver includes the behavior plan, the rationale behind the decision, a route replanning flag, and the designated route starting point (S3.3), as shown in Fig 2.
@@ III-B Integration with Autonomous Vehicle System
We design StuckSolver as an add-on module that seamlessly integrates into existing AV systems to mitigate the inference latency challenges of LLMs. Implemented as a high-level reasoning module, StuckSolver communicates with the AV’s primary modules via structured APIs. It remains non-intrusive to the AV’s default behavior—only intervening when a stuck situation is detected. To illustrate its integration into an AV system, we use the Behavior Agent in CARLA  [17]. StuckSolver interacts with the AV system through the following key connections:
Vehicle Perception-StuckSolver. The original Behavior Agent has no perception module, so we add one with camera, Lidar, Inertial Measurement Unit (IMU), Global Navigation Satellite System (GNSS) sensors to capture surrounding objects and ego state data. The vehicle perception module directly transmits the raw image to StuckSolver, enabling it to analyze the environment ahead and extract key information. For the object measurements, they are obtained from the Lidar with the CenterPoint algorithm  [39]. Before being passed to StuckSolver, both the object measurements and the ego state data are converted into string format and then provided as input.
StuckSolver-Vehicle Route Planning. When route replanning is necessary, StuckSolver determines suitable waypoints on specific lanes as potential starting points using road map data offered by the vehicle’s route planning module. It will output the route replanning? and a route starting point to the route planning module for execution.
StuckSolver-Vehicle Decision making. If the AV is stuck, StuckSolver outputs the behavior plan as a list, replacing the AV’s original decision output. All behaviors included in the behavior plan are then sequentially executed by the corresponding control algorithms within the AV’s control module. If the vehicle is not stuck, StuckSolver simply returns None, which will not affect the AV’s original decision.
Vehicle Control Module. The control module is responsible for executing the planned trajectory by generating low-level control commands, such as steering angle, throttle, and braking. We enrich the module with algorithms for different behavior. Lane keeping is controlled using PID algorithms for lateral steering. Car-following relies on the Intelligent Driver Model (IDM)  [38] for longitudinal control. Lane changing is achieved by using separate PID controllers for both lateral and longitudinal directions.

### v2

@@ Abstract
Despite significant advancements in recent decades, autonomous vehicles (AVs) continue to face challenges in navigating certain traffic situations that human drivers handle with ease. In such cases, AVs may become immobilized, leading to disruptions in overall traffic flow. Existing recovery solutions, including remote intervention, which is costly and inefficient, and manual takeover, which excludes non-drivers and limits accessibility, remain inadequate. This paper introduces StuckSolver, a novel Large Language Model (LLM) driven recovery framework that enables AVs to resolve immobilization scenarios through self-reasoning and passenger-guided decision making. StuckSolver is designed as a plug-in add-on module that operates on top of the AV’s existing perception, planning, and control stack without requiring modifications to its internal architecture. It interfaces directly with standard sensor data streams to detect immobilization states, interpret the environmental context, and generate high-level recovery commands that can be executed by the AV’s native planner. StuckSolver is integrated with CARLA’s rule-based Behavior Agent and evaluated using the Bench2Drive benchmark. Experimental results show that StuckSolver achieves driving scores and success rates on par with state-of-the-art end-to-end approaches through self-reasoning alone, and exhibits additional improvements when passenger guidance is incorporated. This confirms that StuckSolver provides robust operational performance within a lightweight and interpretable framework.
@@ I Introduction
After years of development, AVs have progressed beyond controlled environments such as airports and ports. Today, they are increasingly navigating more complex urban environments. Companies like Baidu and Waymo have deployed large fleets of robotaxis in cities such as Beijing, Wuhan, Phoenix, San Francisco, and Los Angeles, providing public services. For instance, Waymo’s robotaxis offer approximately 150,000 paid rides weekly, collectively traveling over one million miles  [5]. This expansion marks a significant step toward integrating AVs into everyday urban mobility.
@@
Figure 1: Illustration of AV immobilization in traffic. The white AV, carrying non-driving passengers, travels in the right lane. When the leading blue vehicle becomes disabled and the left lane is obstructed by a paper bag, the AV’s re-planning module fails to generate an alternative trajectory due to the perceived obstruction. Although the bag is a benign and traversable object, the AV misclassifies it as a non-traversable obstacle, resulting in a complete stop and prolonged immobilization.
Figure 2: An overview of the LLM-assisted AV recovery method.
@@
Although AVs can complete most trips, they still face significant challenges in navigating certain traffic scenarios. Factors such as insufficient semantic understanding of surrounding entities and dense interactions at critical spatial locations can cause the decision-making or planning module to fail to generate, or mistakenly filter out, valid solutions, ultimately resulting in immobilization (i.e., becoming stationary or unable to operate normally)  [36, 37]. More critically, when this occurs, they frequently lack effective recovery mechanisms to resume normal operations, as illustrated in Fig. 1. Such incidents can cause substantial traffic disruptions and inconvenience for passengers. For example, a Waymo robotaxi in California became trapped in a roundabout after failing to generate a feasible exit trajectory. The vehicle circled 17 times before being remotely rescued by engineers  [6]. Passenger intervention is another recovery method. For instance, Tesla’s Full Self-Driving (FSD) system exits autonomous mode upon detecting operational limits and prompts the driver to take control [11]. However, these existing methods have notable limitations. A remote control requires significant financial and human resources to ensure prompt assistance. Passenger intervention, on the other hand, relies on the occupant’s ability to drive. This leaves elderly, disabled, or non-driving passengers without a viable solution. These challenges underscore the need for more robust and inclusive approaches to enhance AV resilience and accessibility.
@@
LLMs offer a promising solution due to their extensive knowledge and advanced reasoning capabilities [12]. On one hand, LLMs have the potential to understand complex traffic scenarios and generate logical driving decisions to rescue immobilized AVs, eliminating the need for costly remote control resources [15]. On the other hand, LLMs can process natural language, enabling them to act as a bridge for passenger intervention [16]. By interpreting human commands, LLMs can advise appropriate AV behaviors, making AV rides more accessible and equitable. This capability is particularly beneficial for individuals who do not or cannot drive, ensuring that AV technology serves a broader and more inclusive population.
@@
In this paper, we propose StuckSolver, an LLM-based recovery method designed to address AV immobilization in traffic. StuckSolver is developed as an add-on module that can be seamlessly integrated into existing rule-based AV systems, requiring minimal adaptation. It monitors the surrounding traffic conditions and the AV’s operational status. When immobilization is detected, StuckSolver intervenes by generating recovery decisions and overriding the AV’s original operational plans. In scenarios where StuckSolver is unable to resolve the immobilization autonomously, passengers can provide guidance, with StuckSolver acting as an intermediary to interpret and execute human commands. To evaluate its efficacy, StuckSolver is evaluated on the corner-case-rich benchmark Bench2Drive  [35]. Simulation results demonstrate that StuckSolver effectively recovers immobilized AVs, either by leveraging its own knowledge and reasoning capabilities or by interpreting passenger instructions. These findings highlight the potential of StuckSolver to enhance AV resilience in complex traffic scenarios.
@@ II Related work
However, most existing research focuses on isolated tasks (such as perception, decision-making, or planning), letting LLMs operate independently and evaluating their performance within only a single module. While this can demonstrate improvements to specific components of the AV system, it offers little clear evidence of the enhancement in overall driving capability. Some studies have integrated LLMs into a complete AV system [33, 34] and evaluated the driving performance of LLM-powered AVs. However, those efforts primarily focus on normal driving conditions and largely neglect edge cases. Consequently, AVs may struggle in complex or unforeseen scenarios, increasing the risk of prolonged immobilization and compromising both safety and efficiency. Moreover, most of the proposed methods rarely allow passenger intervention, such as providing necessary information or decision guidance. This lack of communication misses the opportunity to further enhance the AV’s ability to navigate complex traffic scenarios.
@@ III LLM-powered Recovery Method
In this section, we present an LLM-powered recovery method designed to enable AVs to autonomously escape from immobilized situations or recover under passenger guidance through natural language commands, eliminating the need for direct driving intervention. The overall framework is illustrated in Fig. 2.
@@ III-A StuckSolver
Our approach transforms GPT-4o [27] into an intelligent agent called StuckSolver by leveraging prompt engineering [28], Chain-of-Thought (CoT) reasoning [29], and OpenAI’s Function Calling API [30]. Operating in a zero-shot mode without any task-specific fine-tuning or additional training, StuckSolver is integrated into the AV system as an add-on module. Its primary role, defined in the system prompt, is to help the AV recover from immobilization or follow passenger guidance to facilitate recovery. When immobilization is detected, StuckSolver initiates a structured CoT reasoning process that sequentially performs environment observation, multi-step analysis, and decision-making—ultimately generating a recovery plan that includes a route replanning flag, route start point, and a detailed action plan to assist the vehicle in regaining normal operation, as illustrated in Fig. 3.
@@
Figure 3: Chain of Thought for StuckSolver. S is an abbreviation for Step. {immobilized? reason} represents whether the AV is immobilized and the corresponding reason.
TABLE I: Results on Bench2Drive Closed-loop Simulations
Method Scheme Modality Closed-loop Metrics
DS $\uparrow$ SR(%) $\uparrow$ Efficiency $\uparrow$ Comfort $\uparrow$
TCP-traj Imitation learning Image 59.90 30.00 76.54 18.08
AD-MLP Imitation learning State 18.05 0.00 48.45 22.63
UniAD-Base Imitation learning Image 45.81 16.36 129.21 43.58
VAD Imitation learning Image 42.35 15.00 157.94 46.01
ThinkTwice Imitation learning Image 62.44 31.23 69.33 16.22
DriveAdapter Imitation learning Image 64.23 30.22 70.22 16.01
GenAD Imitation learning Image 44.81 15.90
DriveTrans Imitation learning Image 63.46 35.01 100.64 20.78
MomAD Imitation learning Image 44.54 16.71 170.21 48.63
Raw2Drive (SOTA) Reinforcement learning Image 71.36 50.24 214.17 22.42
Behavior Agent Rule-based Image, LiDAR 48.70 18.21 68.37 43.68
Behavior Agent + StuckSolver (Ours) Rule-based+Zero-shot Image, LiDAR 65.23 36.32 87.49 46.74
Behavior Agent + StuckSolver (Ours) Rule-based+Zero-shot Image, LiDAR 70.89 50.01 87.78 47.31
denotes expert feature distillation. indicates the use of human guidance in certain scenarios.
The Behavior Agent in CARLA does not include a perception module, which we add by integrating camera and LiDAR sensors.
@@
Observation. Observation is the first and most fundamental step in StuckSolver’s reasoning process. A thorough observation phase ensures that the system collects all necessary information before advancing to analysis and decision-making. StuckSolver captures raw images from the vehicle’s front-view camera and extracts semantic information, including traffic control factors and traffic participants. Traffic control factors ($TC$), such as traffic lights state ($TC_{tl}$), traffic signs content ($TC_{ts}$), and work zones ($TC_{wz}$). Traffic participants ($TP$), including vehicles ($TP_{v}$), pedestrians ($TP_{p}$), and obstacles ($TP_{o}$). For $TP_{v}$ and $TP_{p}$, they include type, lane position (ego lane or adjacent lane), and intent. In addition to semantic interpretation, StuckSolver augments semantic insights with quantitative measurements (e.g., distance, velocity) derived from the perception module, enabling numerically grounded reasoning and subsequent analysis.
@@
Analysis. At this stage, the goal is to determine whether the vehicle is stuck and, if so, identify the cause. The analysis process follows two possible branches: with or without passenger guidance. When passenger guidance is available (S2.1), StuckSolver interprets the passenger’s intent based on observations and formulates a behavior plan to assist the immobilized vehicle in resuming movement. In the absence of passenger guidance (S2.2-2.5), StuckSolver first checks the destination flag to determine whether it has stopped en route to its destination. If the vehicle is still in motion, StuckSolver outputs None, allowing autonomous operation to continue without intervention. However, if the vehicle has not yet reached its destination and its speed remains below the minimum threshold $v_{min}$ (set to 5km/h in this study) for longer than the stationary duration $\triangle t$ (set to 1 second in this study), it is flagged as potentially immobilized, triggering further analysis. StuckSolver begins by assessing whether any traffic control elements are influencing the vehicle’s movement. It checks $TC$ to identify possible causes of immobilization. Also, StuckSolver evaluates, based on $TP$, whether the AV has stopped to yield to surrounding traffic participants, or to avoid a potential collision with obstacles. Finally, it provides the analysis results—indicating whether the vehicle is immobilized and the underlying cause—to the decision-making process.
@@
Decision-making. If a passenger-suggested behavior plan is available (S3.1), StuckSolver evaluates whether route replanning is necessary to reach the destination. In the absence of passenger guidance (S3.2), StuckSolver assesses the ego state, map information, and available behavior options to generate a safe and traffic-compliant behavior plan. It also determines whether route replanning is required and identifies the appropriate starting point for the new route. Finally, the structured output of StuckSolver includes the behavior plan, the rationale behind the decision, a route replanning flag, and the designated route starting point (S3.3).
@@ III-B Integration with Autonomous Vehicle System
We design StuckSolver as an add-on module that seamlessly integrates into existing AV systems to mitigate the inference latency challenges of LLMs. Implemented as a high-level reasoning module, StuckSolver communicates with the AV’s primary modules via structured APIs. It remains non-intrusive to the AV’s default behavior, intervening only when a stuck situation is detected. Its integration is demonstrated using the Behavior Agent in CARLA  [17]. The data flow between StuckSolver and the AV system is as follows:
Data Flow from the AV System to StuckSolver. The original Behavior Agent has no perception module, so we add a camera, LiDAR, Inertial Measurement Unit (IMU), and Global Navigation Satellite System (GNSS) sensors to capture surrounding objects and ego state data. The vehicle perception module directly transmits the raw image to StuckSolver, enabling it to analyze the environment ahead and extract key information. For the object measurements, they are obtained from the LiDAR with the CenterPoint algorithm  [39]. Before being passed to StuckSolver, both the object measurements and the ego state data are converted into string format and then provided as input.
In addition, the route planning module provides StuckSolver with map information, such as the IDs of the current and adjacent lanes, as well as the candidate waypoints available for route replanning. The motion planning module, on the other hand, outputs a list of currently available driving behaviors (e.g., car following, lane keeping, and lane changing ) that StuckSolver can consider when generating its recovery plan.
Figure 4: Experimental results for three representative scenarios: pedestrian crossing, unexpected door opening, and cyclist following. The left panels display RGB images captured by the ego vehicle’s front-view camera, while the right panels illustrate the corresponding decision-making processes of the StuckSolver.
Data Flow from StuckSolver to the AV System. After decision-making, if the vehicle is not stuck, StuckSolver simply returns None, which will not affect the AV’s original decision. If the AV is stuck, StuckSolver outputs the behavior plan as a list, replacing the AV’s original decision output. All behaviors included in the behavior plan are then sequentially executed by the corresponding control algorithms within the AV’s control module. When route replanning is necessary, StuckSolver determines suitable waypoints on specific lanes as potential starting points. It will output a replanning flag, route replanning?, and a route starting point to the route planning module for execution.